the supply and demand for broccoli are described by the following equations;supply QS=6p-80 Demand QD= 120-4p
a. grapgh the suppply curve and the demand curve .what is the equalibrium price and quantity?
b.calculate consumer surplus,producer surplus and total surplus at the equalibrium.
c.if a dictator who hated brocooli were to ban the vegetables,who would bear the larger burden_the buyers or sellers of brocooli?
1
Expert's answer
2020-05-12T10:44:07-0400

a.





"Q_D=Q_S"

"120-4p=6p-80"


"10p=200"


"p=20"


"Q=40"

b.

consumer surplus is "\frac {1}{2}(30-20)\times 40=200"


producer surplus is "\frac{1}{2}(20-\frac{80}{6})\times40=133"


 total surplus  is 200+133=333


c. If the dictator had banned broccoli, sellers would have incurred large losses, since their losses would have included the cost of broccoli cultivation and lost profits.